OpenStudy (anonymous):

how do i find the mass of an object with 3m/s

OpenStudy (anonymous):

Only velocity is given??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is the question "A children's roller coaster is released from the top of a track, if its maximum speed at ground level is 3 m/s. find the height that its released from.

OpenStudy (raffle_snaffle):

kinetic energy

OpenStudy (mrnood):

you do not need the mass At the 'ground level' it has a speed of 3 m/s which gives it a KE of 0.5 mv^2 At the top of the track it has a PE of mgh The PE is converted to PE when it rolls down so: mgh = 0.5mv^2 and you can see that the mass cancels out h=0.5v^2/g

OpenStudy (anonymous):

@jose8787 u=0 m/s v=3 m/s a=g=9.81 m/s^2 \[v ^{2}=u ^{2}+2as \] \[s=v ^{2}/2g\] s=h= 9/s*9.81 s=.45 m